OpenStudy (itsmichelle29):

Simplify root(1-sintheta)(1+sintheta)

OpenStudy (anonymous):

\[\sqrt{(1-sin{\theta})(1+sin{\theta})}\] ?

OpenStudy (anonymous):

Is that what you meant?

OpenStudy (itsmichelle29):

Yup

OpenStudy (itsmichelle29):

±sin θ cos θ ±tan θ square root sine theta

OpenStudy (itsmichelle29):

those are the answerscan u help me

OpenStudy (anonymous):

\(\sqrt{(1-sin{\theta})(1+sin{\theta})}\) \(\sqrt{(1-sin^2{\theta })}\) But \(cos^2{\theta} + sin^2{\theta} = 1\), then \(\sqrt{cos^2{\theta}}\) \(\pm cos{\theta}\)

OpenStudy (itsmichelle29):

so would the answer be costheta only

OpenStudy (anonymous):

yes

OpenStudy (itsmichelle29):

omg thank you can u help me with one more only

OpenStudy (anonymous):

sure^^

OpenStudy (itsmichelle29):

@M4thM1nd

OpenStudy (anonymous):

\(cos(x-\pi/2) = sin(x)\) and \(sin(x - \pi/2)=-cos(x)\)

OpenStudy (anonymous):

So, which one you think is the right answer?

OpenStudy (itsmichelle29):

1st one

OpenStudy (itsmichelle29):

i think ... idk

OpenStudy (anonymous):

No. From the graph we see that f(x) at x = 0 is 0, that means f(x) is some function of sin(x), since sin(x) at x = 0 is 0

OpenStudy (itsmichelle29):

so the answer is a

OpenStudy (anonymous):

In that case, the first one we have \(sin(x-\pi/2) = -cos(x)\), so that can't be the right answer

OpenStudy (itsmichelle29):

okay the right answer is ...

OpenStudy (anonymous):

If we check the last option, we have \(cos(x-\pi/2) = sin(x)\), but from the graph we also get the information that f(x) at x = pi/2 is equal to -4. But sin(x) at x = pi/2 is equal to 1. That tells us that we need a -4 multiplying this sin(x). So... The last option is the correct answer

OpenStudy (anonymous):

do you understand?